# Tomographic redshift dipole: Testing the cosmological principle

### v3

@@ 2 The redshift dipole
The clustering dipole’s main contribution comes from peculiar motions, and can be estimated by computing the bulk flow velocity, which is the expected root mean square velocity of all sources in a survey. We found that the clustering dipole is negligible for the dataset we consider, having a magnitude smaller than 2% of the expected $\beta$ [31]. Therefore, from now on, we will simply adopt $\boldsymbol{\Delta_{1}}=-\boldsymbol{\beta}$, i.e., we disregard terms of the order $\mathcal{O}(\beta^{2})$.
@@ 3 Data
Figure 2: Redshift distribution of SDSS data used. The bin size adopted is $\Delta z=0.001$, the same used in the estimator. Dashed lines indicate data constraints. From left to right, the first dashed line indicates the upper bound for CMASS eBOSS data, $z<0.75$, implemented to prevent estimator misbehavior with sparsely populated bins. Similarly, for LRG data, we adopt $z<0.9$, denoted by the second dashed line.
@@
The Emission Line Galaxy (ELG) eBOSS catalog was not considered as it covers merely $\sim$3of the sky. Given our aim to assess the largest scale modulation, the extent of sky coverage is crucial; hence, incorporating the ELG catalog yields negligible enhancement to our analysis. BOSS objects with $z<0.4$ (mostly LOWZ galaxies) were not considered because it was not possible to minimize the estimator. We will come back to this issue in Section 5. Furthermore, CMASS eBOSS objects with $z>0.75$ were excluded, leading to a 4.6% reduction in the original catalog, and LRG eBOSS objects with $z>0.9$ were also omitted, accounting for a 6.6% reduction. These minor exclusions were necessary to prevent estimator misbehavior, as the estimator’s minimization becomes unstable in sparsely populated redshift bins (we adopted thin bins of $\Delta z=0.001$).
@@ 4.1 Estimator
accounting for redshift failures ($w_{i,\mathrm{noz}}$), fiber collisions ($w_{i,\mathrm{fc}}$), and imaging systematics ($w_{i,\mathrm{sys}}$). Spectroscopic redshift measurement errors are negligible ($\sim 10^{-4}$[40, 33, 41] and thus not considered.
@@
Figure 4: Estimator diagram. Diagram of the $1+z$ spherical shell bin, its monopole and data. The estimator minimizes the difference between the observed $1+z_{i}$ (black points) and the monopole $\Delta^{dD}_{0,{\rm bin}}$ (the average of $1+z_{i}$ in the rest frame, described by the red line) times a dipole modulation $\delta(-\boldsymbol{\Delta}_{1},\boldsymbol{\hat{n}}^{\prime}_{i})$. The direction vector $\boldsymbol{\hat{n}}^{\prime}_{i}$ originates from the observer at the center of the shell.
Figure 5: Distribution of monopoles ($\Delta_{0}$, average redshift) for the North (NGC) and South (SGC) Galactic Caps mocks. $\bar{\Delta}^{\rm NGC}_{0}$ represents the average NGC monopole, i.e., average $1+z$. It is clear that the hemispheres have different depths.
@@ 4.2 Bias estimation and removal
To construct the bias map for both the eBOSS and BOSS surveys, we first Doppler boosted 100 mocks, randomly selected from the original pool of 1000 for each dataset, across 12 distinct directions identified by the pixel centers in the HEALPix $N_{\rm side}=1$ scheme [^1] HEALPix [42, 43], or Hierarchical Equal Area isoLatitude Pixelization, is an efficient system widely used in astrophysics and cosmology for pixelating the celestial sphere, enhancing the analysis and visualization of astronomical datasets.. These mocks were simulated at a velocity of 370 km/s, consistent with the kinematic interpretation of the CMB dipole. Our tests confirm that the bias map remains reliable across various dipole magnitudes. Additionally, we validated the accuracy of our estimator using the 4000 original mocks (1000 per catalog) without Doppler effects, successfully achieving expected results with bias correction, as depicted in Fig. 6.
@@ 5 Results and Discussion
Figure 9: Final dipole measurements. Mollweide projection displaying the dipole measurements in galactic coordinates, with contours representing the $1\sigma$ and $2\sigma$ confidence regions for combined measurements. The star symbol denotes the position of the CMB dipole for reference.

### v4

@@ 2 The redshift dipole
The clustering dipole’s main contribution comes from peculiar motions, and can be estimated by computing the bulk flow velocity, which is the expected root mean square velocity of all sources in a survey (see equation D.2 from [31] for the exact definition). We found that the clustering dipole is negligible for the dataset we consider, with a magnitude less than 2% of the expected $\beta$ [31]. Consequently, we will adopt $\boldsymbol{\Delta_{1}}=-\boldsymbol{\beta}$, thereby disregarding terms of the order $\mathcal{O}(\beta^{2})$. In our approximation, where the only source of the dipole is the velocity, this implies a unique definition of $\boldsymbol{\beta}$ as $\boldsymbol{\beta}=-\boldsymbol{\Delta_{1}}$, ensuring that the dipole is null in the rest frame.
@@ 3 Data
Figure 2: Redshift distribution of SDSS data used. The bin size adopted in the plot is $\Delta z=0.001$. Dashed lines indicate data constraints. From left to right, the first dashed line indicates the upper bound for CMASS eBOSS data, $z<0.75$, implemented to prevent estimator misbehavior with sparsely populated bins. Similarly, for LRG data, we adopt $z<0.9$, denoted by the second dashed line.
@@
The Emission Line Galaxy (ELG) eBOSS catalog was not considered as it covers merely $\sim$3of the sky. Given our aim to assess the largest scale modulation, the extent of sky coverage is crucial; hence, incorporating the ELG catalog yields negligible enhancement to our analysis. BOSS objects with $z<0.4$ (mostly LOWZ galaxies) were not considered because it was not possible to minimize the estimator. We will come back to this issue in Section 5. Furthermore, CMASS eBOSS objects with $z>0.75$ were excluded, leading to a 4.6% reduction in the original catalog, and LRG eBOSS objects with $z>0.9$ were also omitted, accounting for a 6.6% reduction. These minor exclusions were necessary to prevent estimator misbehavior, as the estimator’s minimization becomes unstable in sparsely populated redshift bins.
@@ 4.1 Estimator
accounting for redshift failures ($w_{i,\mathrm{noz}}$), fiber collisions ($w_{i,\mathrm{fc}}$), and imaging systematics ($w_{i,\mathrm{sys}}$).
The spectroscopic redshift measurement errors are approximately $2\times 10^{-4}$ for LRG eBOSS, CMASS eBOSS, and BOSS $(z\geq 0.4)$, and around $5\times 10^{-4}$ for QSO eBOSS objects [40, 33, 41], as shown in Fig. 4. The 90th percentiles of the error distributions for QSO eBOSS, LRG eBOSS, CMASS eBOSS, and BOSS $(z\geq 0.4)$ are 0.0011, 0.00025, 0.00037, and 0.00020, respectively. Consequently, we conservatively choose redshift bin widths that are approximately three times these values, adopting $\Delta z=0.001$ for LRG eBOSS, CMASS eBOSS, and BOSS $(z\geq 0.4)$, and $\Delta z=0.003$ for QSO eBOSS. As verified in Appendix C, the error evolution for the QSO eBOSS sample across different bins remains consistent with the chosen bin size throughout the entire redshift range.
Figure 4: Redshift error distribution for each sample. The median error from left to right is 0.00052, 0.00017, 0.00025, 0.00013.
@@
Figure 5: Estimator diagram. Diagram of the $1+z$ spherical shell bin, its monopole and data.The dotted curve represents the observed 1+z for several objects (dots). The estimator minimizes the difference between the observed $1+z_{i}$ (black points) and the monopole $\Delta^{dD}_{0,{\rm bin}}$ (the average of $1+z_{i}$ in the rest frame, described by the red line) times a dipole modulation $\delta(-\boldsymbol{\Delta}_{1},\boldsymbol{\hat{n}}^{\prime}_{i})$. The direction vector $\boldsymbol{\hat{n}}^{\prime}_{i}$ originates from the observer at the center of the shell.
Figure 6: Distribution of monopoles ($\Delta_{0}$, average redshift) for the North (NGC) and South (SGC) Galactic Caps mocks. $\bar{\Delta}^{\rm NGC}_{0}$ represents the average NGC monopole, i.e., average $1+z$. It is clear that the hemispheres have different depths in the QSO eBOSS and LRG eBOSS samples. Even for CMASS eBOSS and BOSS $(z\geq 0.4)$ the difference is not insignificant for our estimator, as it provides a systematic shift of $\sim 3.5\times 10^{-4}$, while the expected modulation is $\sim 10^{-3}$.
@@ 4.2 Bias estimation and removal
To construct the bias map for both the eBOSS and BOSS surveys, we first Doppler boosted 100 mocks, randomly selected from the original pool of 1000 for each dataset, across 12 distinct directions identified by the pixel centers in the HEALPix $N_{\rm side}=1$ scheme [^1] HEALPix [42, 43], or Hierarchical Equal Area isoLatitude Pixelization, is an efficient system widely used in astrophysics and cosmology for pixelating the celestial sphere, enhancing the analysis and visualization of astronomical datasets.. These mocks were simulated at a velocity of 370 km/s, by multiplying the $1+z$ value of each object by the expected Doppler factor as defined by equation (2.2), consistent with the kinematic interpretation of the CMB dipole. Our tests confirm that the bias map remains reliable across various dipole magnitudes. Additionally, we validated the accuracy of our estimator using the 4000 original mocks (1000 per catalog) without Doppler effects, successfully achieving expected results with bias correction, as depicted in Fig. 7.
@@ 5 Results and Discussion
Figure 10: Final dipole measurements. Mollweide projection displaying the $\boldsymbol{\beta}$ measurements in galactic coordinates, representing the negative of the redshift dipole ($\boldsymbol{\Delta}_{1}=-\boldsymbol{\beta}$). Contours indicate the $1\sigma$ and $2\sigma$ confidence regions for the combined measurements. The star symbol denotes the position of the CMB dipole for reference ($\boldsymbol{\Delta}_{1,\rm{CMB}}=\boldsymbol{\beta}$).
@@ Appendix B Fisher’s method
Appendix C Redshift error evolution
We examined in Fig. 13 the potential evolution of redshift errors with increasing redshift, confirming that a bin width of 0.003 is sufficient across the entire redshift range.
Figure 13: Redshift error evolution of the QSO eBOSS sample. Dashed lines mark the median error.